# Data-Driven Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

### v1

@@ 4 Derivation of GSA indices from PDP and ICE
$I_{pdp}\left(x_{j}\right)=\sqrt{\text{ Var}_{x_{j}}[\hat{f}_{pdp}(x_{j})]}\approx\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\hat{f}_{pdp}\left(x_{j}^{(k)}\right)-\frac{1}{K}\sum_{k=1}^{K}\hat{f}_{pdp}\left(x_{j}^{(k)}\right)\right)^{2}},$ (9)
where $K$ is the number of samples drawn from $\rho_{x_{j}}(x_{j})$.
$I_{pdp}$ corresponds exactly to the (unnormalised) first‑order effect in the Hoeffding–Sobol (functional ANOVA) decomposition [13]. However, it should not be interpreted as a proportion of total variance unless normalised by the total variance itself, because it only measures the spread of the conditional mean. Moreover, since PDP is computed by averaging model predictions over the complementary variables, its curve can blur interaction effects. This is precisely why ICE curves (and their dispersion) are needed when heterogeneity or interaction is suspected [18, 38].
A similar metric for assessing the interaction term can also be computed. Formally, let $\boldsymbol{x}_{S}=\{x_{i},x_{j}\}$ be the set of two variables that we have an interest in. First, build the joint partial dependence function $\hat{f}_{pdp}(x_{i},x_{j})$. From $\hat{f}_{pdp}(x_{i},x_{j})$, we can calculate $I_{pdp}(x_{i}|x_{j})$ for each value of $x_{j}$, in other words, we compute the partial dependence function of $x_{i}$ on the joint dependence function $\hat{f}_{pdp}(x_{i},x_{j})$. The standard deviation of $I_{pdp}(x_{i}|x_{j})$ is then computed, and the same is done for $I_{pdp}(x_{j},x_{i})$. The standard deviation values for both $I_{pdp}(x_{i}|x_{j})$ and $I_{pdp}(x_{j}|x_{i})$ are then averaged to yield the interaction effect. Regardless, Eq. (9) still represents only the average effect and might cause misunderstanding, which is why we introduce ICE-based feature importance, as explained in the next section.
@@ 4.1 ICE-based feature importance
In this paper, we address the limitations of PDP-based feature importance by introducing an ICE-based feature importance metric. While conceptually similar to PDP-based importance, the key distinction is that we first quantify the feature importance for each individual ICE curve before aggregating. Formally, the variance of a single ICE curve with respect to $x_{j}$, given the $i$-th realization of the complementary features in $\Omega_{C}$ (denoted as $\boldsymbol{x}_{C}^{(i)}$), is calculated as follows:
@@
$\mu_{I_{\mathrm{ice}},x_{j}}=\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\int_{\Omega_{C}}I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{c}\right)\rho_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{c})d\boldsymbol{x}_{c},$ (14)
@@
$\sigma_{I_{\mathrm{ice}},x_{j}}^{2}=\mathrm{Var}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\int_{\Omega_{C}}\left(I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{c})-\mu_{I_{\mathrm{ice}},x_{j}}\right)^{2}\rho_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{c})d\boldsymbol{x}_{c}.$ (15)
@@
These two metrics serve distinct but complementary purposes. The first metric, $\mu_{I_{\mathrm{ice}},x_{j}}$, measures global importance. Unlike PDP-based importance, which averages the predictions (often leading to cancellation effects), $\mu_{I_{\mathrm{ice}},x_{j}}$ averages the magnitude of variations, thereby mitigating the issue of heterogeneous effects canceling each other out. Conversely, the second metric, $\sigma_{I_{\mathrm{ice}},x_{j}}$, assesses the influence of interactions. A high dispersion in $I_{\mathrm{ice}}$ indicates that the impact of $x_{j}$ varies significantly depending on the values of the complementary features $\boldsymbol{x}_{C}$, which is a hallmark of strong feature interactions. These metrics can be visualized simultaneously; for example, using a bar plot for $\mu_{I_{\mathrm{ice}},x_{j}}$ with $\sigma_{I_{\mathrm{ice}},x_{j}}$ represented as error bars.
@@
Equation (10) defines the conditional variance of the ICE curve for instance $\boldsymbol{x}_{C}^{(i)}$, and Eq. (12) is its empirical square root. When we then average and compute the spread of $I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C}^{(i)})$ across instances, we obtain interpretable quantities in the same units as the model output. One could instead use the conditional variance itself, without taking a square root, as the instance-wise importance measure. That choice has a formal appeal: by the law of total variance, $\mathrm{Var}(Y)=\mathrm{Var}_{X_{j}}\bigl(\mathbb{E}[Y\mid X_{j}]\bigr)+\mathbb{E}_{X_{j}}\bigl[\mathrm{Var}(Y\mid X_{j})\bigr],$ and in classical Sobol’ analysis (assuming input independence), one identifies the first-order component with $\mathrm{Var}_{X_{j}}(\mathbb{E}[Y\mid X_{j}])$ and the complementary term $\mathbb{E}[\mathrm{Var}(Y\mid X_{j})]$ as the residual / interaction-plus component. This decomposition underlies how Sobol’ indices are defined [41]. By contrast, standard deviations are more interpretable (same units as $Y$), but do not decompose additively in the same way. Thus, the two metrics are complementary, and we recommend using the variance for variance decomposition and additivity, while the standard deviation is of better use for interpretability in the original unit of model output.
Recent work has established a rigorous connection between conditional variability summaries and optimal transport-based GSA [42, 43]. Specifically, in [43], the authors propose a sensitivity index based on the squared 2-Wasserstein distance ($W_{2}^{2}$) to quantify the discrepancy between the unconditional model output distribution $\mathcal{L}(Y)$ and the conditional distribution $\mathcal{L}(Y|x_{j})$ [43].
Under the assumption that the distributions are Gaussian or elliptical, this distance admits a closed-form decomposition known as the Wasserstein-Bures semimetric:
@@
$W_{2}^{2}\big(\mathcal{L}(Y),\mathcal{L}(Y|x_{j})\big)=\underbrace{\|\mu_{Y}-\mu_{Y|x_{j}}\|_{2}^{2}}_{\text{Mean Shift}}+\underbrace{\text{Tr}\left(\Sigma_{Y}+\Sigma_{Y|x_{j}}-2(\Sigma_{Y}^{1/2}\Sigma_{Y|x_{j}}\Sigma_{Y}^{1/2})^{1/2}\right)}_{\text{Covariance Shift}}$ (16)
@@
Our proposed ICE-based metrics align with this theoretical framework by explicitly separating the global average effect (related to the mean shift) from the dispersion of individual ICE curves (related to the covariance shift). This ensures that features influencing the model’s variability, even without shifting the global mean, are correctly identified as important. The key observation is that the partial dependence function $\hat{f}_{pdp}(x_{j})$ averages out the variability of $\hat{f}(x_{j},\boldsymbol{x}_{C})$ across all values of $\boldsymbol{x}_{C}$, potentially smoothing out variations that would otherwise be captured in the ICE curves. Thus, the variability captured by $I_{ice}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)$ will always account for more localized variations specific to $\boldsymbol{x}_{C}^{(i)}$, which are not captured by the globally averaged $\hat{f}_{pdp}(x_{j})$. As a result, we demonstrate, in a certain setting, that:
$\mathbb{E}_{x_{C}}\left[I_{ice}(x_{j};x_{C})\right]\geq I_{pdp}(x_{j}).$ (17)
More precisely, we demonstrate that the inequality holds for several classes of functions, most notably for those expressible as a finite expansion of a linear combination of orthogonal bases and functions with some combination of additive and multiplicative separability, as detailed in Appendix A. Since the PCE surrogate model itself constitutes a finite expansion of such bases, the inequality is satisfied when PCE is employed.
@@ 4.2 ICE-based correlation values
Beyond quantifying GSA, it is also essential to assess how interactions modify the relationship between inputs and predictions. This enables the identification of variables that change from positive to negative associations (or retain their direction but vary in magnitude) when interacting with others. Such insights are invaluable in design exploration, guiding more informed decisions on controlling design parameters during optimization. To complement the ICE-based feature importance, we introduce the ICE-based correlation values which quantify the correlation between ICE curves and the partial dependence function, written as
@@
where Corr denotes the correlation (we use the Pearson correlation coefficient). That is, we compute the correlation coefficient between the ICE curve $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$ for a specific instance $\boldsymbol{x}_{C}^{(i)}$ and the corresponding partial dependence function $\hat{f}_{pdp}(x_{j}))$ for the variable $x_{j}$.
@@
which is essentially the standard deviation of the ICE correlation values, and can be computed from a Monte Carlo simulation. If $\sigma_{\rho}=0$, this indicates that the partial dependence function is similar to ICE curves (in other words, there is no interaction). Conversely, a high value of $\sigma_{\rho}$ suggests that most of the ICE curves differ significantly from the partial dependence line, indicating a strong interaction effect on the trend. Additionally, showing the boxplot of ICE-based correlation values also helps in the assessment.

### v3

@@ 4 Derivation of GSA indices from PDP and ICE
$I_{pdp}\left(x_{j}\right)=\sqrt{\text{ Var}_{x_{j}}[\hat{f}_{pdp}(x_{j})]}\approx\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\hat{f}_{pdp}\left(x_{j}^{(k)}\right)-\frac{1}{K}\sum_{l=1}^{K}\hat{f}_{pdp}\left(x_{j}^{(l)}\right)\right)^{2}},$ (9)
where $K$ is the number of samples drawn from $\pi_{x_{j}}(x_{j})$.
$I_{pdp}$ corresponds exactly to the (unnormalised) first‑order effect in the Hoeffding–Sobol’ (functional ANOVA) decomposition [51]. However, it should not be interpreted as a proportion of total variance unless normalised by the total variance itself, because it only measures the spread of the conditional mean. Moreover, since PDP is computed by averaging model predictions over the complementary variables, its curve can blur interaction effects. This is precisely why ICE curves (and their dispersion) are needed when heterogeneity or interaction is suspected [17].
A similar metric for assessing the interaction term can also be computed. Formally, let $\boldsymbol{x}_{S}=\{x_{i},x_{j}\}$ be the set of two variables that we have an interest in. First, build the joint partial dependence function $\hat{f}_{pdp}(x_{i},x_{j})$. From $\hat{f}_{pdp}(x_{i},x_{j})$, we can calculate $I_{pdp}(x_{i}|x_{j})$ for each value of $x_{j}$, in other words, we compute the partial dependence function of $x_{i}$ on the joint dependence function $\hat{f}_{pdp}(x_{i},x_{j})$. The standard deviation of $I_{pdp}(x_{i}|x_{j})$ is then computed, and the same is done for $I_{pdp}(x_{j}|x_{i})$. The standard deviation values for both $I_{pdp}(x_{i}|x_{j})$ and $I_{pdp}(x_{j}|x_{i})$ are then averaged to yield the interaction effect. Regardless, Eq. (9) still represents only the average effect and might cause misunderstanding, which is why we introduce ICE-based feature importance, as explained in the next section.
@@ 4.1 ICE-based feature importance
In this paper, we address the limitations of PDP-based feature importance by introducing an ICE-based feature importance metric. While conceptually similar to PDP-based importance, the key distinction is that we first quantify the feature importance for each individual ICE curve before aggregating. Formally, the variance of a single ICE curve with respect to $x_{j}$, given the $i$-th realization of the complementary features in $\boldsymbol{\Omega}_{C}$ (denoted as $\boldsymbol{x}_{C}^{(i)}$), is calculated as follows:
@@
$\mu_{I_{\mathrm{ice}},x_{j}}=\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\int_{\boldsymbol{\Omega}_{C}}I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{c}\right)\boldsymbol{\pi}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{c})d\boldsymbol{x}_{c},$ (14)
@@
$\sigma_{I_{\mathrm{ice}},x_{j}}^{2}=\mathrm{Var}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\int_{\boldsymbol{\Omega}_{C}}\left(I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{c})-\mu_{I_{\mathrm{ice}},x_{j}}\right)^{2}\boldsymbol{\pi}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{c})d\boldsymbol{x}_{c}.$ (15)
In practice, these metrics are calculated using Monte Carlo simulations as follows:
$\mu_{I_{\mathrm{ice}},x_{j}}\;\approx\;\hat{\mu}_{I_{\mathrm{ice}},x_{j}}=\frac{1}{N}\sum_{i=1}^{N}I_{\mathrm{ice}}\!\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right).$ (16)
and
$\sigma_{I_{\mathrm{ice}},x_{j}}^{2}\;\approx\;\hat{\sigma}_{I_{\mathrm{ice}},x_{j}}^{2}=\frac{1}{N-1}\sum_{i=1}^{N}\left(I_{\mathrm{ice}}\!\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)-\hat{\mu}_{I_{\mathrm{ice}},x_{j}}\right)^{2},$ (17)
where $\{\boldsymbol{x}_{C}^{(i)}\}_{i=1}^{N}$ is sampled from $\boldsymbol{\pi}_{\boldsymbol{x}_{C}}$ and $N$ is the size of Monte Carlo samples. The proposed ICE-based feature importance method requires evaluating the predictive model on a grid of $K$ values for the feature of interest $x_{j}$ across $N$ realizations of the complementary inputs $\boldsymbol{x}_{C}$. As a result, the computational cost for a single feature scales as $\mathcal{O}(NK)$ model evaluations (the aggregation itself is negligible), and as $\mathcal{O}(mNK)$ when applied to all $m$ input features. This is identical to the computational cost of computing partial dependence plots, since PDPs are obtained by averaging the same set of ICE curves. To put it into perspective, the computational complexity of total Sobol’ indices is $\mathcal{O}(mN)$ using the standard Saltelli estimator for all inputs [44]. In this regard, the proposed method incurs a higher computational cost compared to total Sobol’ indices in exchange for resolving the conditional response structure rather than solely attributing output variance.
@@
These two metrics serve distinct but complementary purposes. The first metric, $\mu_{I_{\mathrm{ice}},x_{j}}$, measures global importance. Unlike PDP-based importance, which averages the predictions (often leading to cancellation effects), $\mu_{I_{\mathrm{ice}},x_{j}}$ averages the magnitude of variations, thereby mitigating the issue of heterogeneous effects canceling each other out. Conversely, the second metric, $\sigma^{2}_{I_{\mathrm{ice}},x_{j}}$, assesses the influence of interactions. A high dispersion in $I_{\mathrm{ice}}$ indicates that the impact of $x_{j}$ varies significantly depending on the values of the complementary features $\boldsymbol{x}_{C}$, which indicates strong feature interactions. These metrics can be visualized simultaneously using a bar plot for $\mu_{I_{\mathrm{ice}},x_{j}}$ with $\sigma_{I_{\mathrm{ice}},x_{j}}$ represented as error bars.
The key difference between the proposed $\mu_{I_{\mathrm{ice}},x_{j}}$ and the total Sobol’ index lies in what aspect of input influence they quantify. The metric $\mu_{I_{\mathrm{ice}},x_{j}}$ characterizes how the input variable $x_{j}$ influences the model response across different conditional settings by measuring the average conditional effect of $x_{j}$ over all realizations of the remaining inputs. In contrast, the total Sobol’ index quantifies the fraction of the output variance attributable to $x_{j}$, including all interaction effects with other variables. While both metrics account for interactions between $x_{j}$ and the remaining inputs, they do so in fundamentally different ways: $\mu_{I_{\mathrm{ice}},x_{j}}$ focuses on conditional response behavior, whereas the total Sobol’ index focuses on variance decomposition. The two metrics can lead to different variable rankings, as illustrated by the first demonstration in Section 4.4. On the other hand, $\sigma_{I_{\mathrm{ice}},x_{j}}^{2}$ is an additional metric that complements $\mu_{I_{\mathrm{ice}},x_{j}}$ by capturing how heterogeneous this effect is with respect to variations in the remaining inputs. This variability is closely related to interaction effects in the Sobol’ sense. The difference is that $\sigma_{I_{\mathrm{ice}},x_{j}}^{2}$ quantifies the variance of each ICE curve across different conditional settings, while Sobol’ interaction indices quantify the contribution of interaction terms to the overall output variance. As such, $\sigma_{I_{\mathrm{ice}},x_{j}}^{2}$ reflects heterogeneity in conditional response behavior, whereas Sobol’ interaction indices reflect variance attribution within a functional ANOVA decomposition.
@@
Equation (11) defines the mean of the ICE curve for instance $\boldsymbol{x}_{C}^{(i)}$, and Eq. (12) is its standard deviation. When we then average and compute the spread of $I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C}^{(i)})$ across instances, we obtain interpretable quantities in the same units as the model output. One could instead use the conditional variance itself, without taking a square root, as the instance-wise importance measure. That choice has a formal appeal: by the law of total variance, $\mathrm{Var}(y)=\mathrm{Var}_{x_{j}}\bigl(\mathbb{E}[y\mid x_{j}]\bigr)+\mathbb{E}_{x_{j}}\bigl[\mathrm{Var}(y\mid x_{j})\bigr],$ and in classical Sobol’ analysis (assuming input independence), one identifies the first-order component with $\mathrm{Var}_{x_{j}}(\mathbb{E}[y\mid x_{j}])$ and the complementary term $\mathbb{E}[\mathrm{Var}(y\mid x_{j})]$ as the residual / interaction-plus component. This decomposition underlies how Sobol’ indices are defined [21]. By contrast, standard deviations are more interpretable (same units as $y$), but do not decompose additively in the same way. Thus, the two metrics are complementary, and we recommend using the variance for variance decomposition and additivity, while the standard deviation is of better use for interpretability in the original unit of model output.
The pseudocode for the calculation of ICE-based feature importance is shown in Algorithm 1. Note that the calculation of the mean ICE-based importance $\mu_{I_{\mathrm{ice}},x_{j}}$ and variance of ICE-based importance $I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)$ are done by using Monte Carlo simulation.
Algorithm 1 ICE-based Feature Importance for a variable $x_{j}$
Trained model $\hat{f}(\boldsymbol{x})$, feature of interest $x_{j}$, $K$ grid points $\{x_{j}^{(k)}\}_{k=1}^{K}$
$\mu_{I_{\mathrm{ice}},x_{j}}$, $\sigma^{2}_{I_{\mathrm{ice}},x_{j}}$
Sample $\{\boldsymbol{x}_{C}^{(i)}\}_{i=1}^{N}$ from $\boldsymbol{\pi}_{\boldsymbol{x}_{C}}$
for $i=1,\dots,N$ do
  Evaluate the $i$-th ICE curve $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$
  Estimate the standard deviation of the $i$-th ICE curve $I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)$ using Eq. (13)
end for
Estimate mean ICE-based importance $\mu_{I_{\mathrm{ice}},x_{j}}$ using Eq. (14)
Estimate variance of ICE-based importance $\sigma^{2}_{I_{\mathrm{ice}},x_{j}}$ using Eq. (15)
Recent work has established a rigorous connection between conditional variability summaries and optimal-transport–based GSA [8, 10]. Here we treat $x_{j}$ as the feature and $x_{j}^{(i)}$ as a specific sampled value of that feature. Denote the unconditional and conditional output distributions by $\mathcal{L}(y)$ and $\mathcal{L}(y\mid x_{j}=x_{j}^{(i)})$, respectively. Define the unconditional mean and covariance as $\mu_{y}=\mathbb{E}[y]$ and $\Sigma_{y}=\mathrm{Var}(y)$, respectively. Also, denote the conditional mean and covariance (pointwise in $x_{j}$) as $\mu_{y\mid x_{j}}=\mathbb{E}[y\mid x_{j}]$ and $\Sigma_{y\mid x_{j}}=\mathrm{Var}(y\mid x_{j})$, respectively. Chiani et al. [10] propose a sensitivity index based on the squared 2–Wasserstein distance, $W_{2}^{2}$, between $\mathcal{L}(y)$ and $\mathcal{L}(y\mid x_{j})$. Under Gaussian or elliptical assumptions, this distance admits the closed-form Wasserstein–Bures decomposition
@@
$W_{2}^{2}\big(\mathcal{L}(y),\mathcal{L}(y\mid x_{j})\big)=\underbrace{\|\mu_{y}-\mu_{y\mid x_{j}}\|_{2}^{2}}_{\text{mean shift}}+\underbrace{\mathrm{Tr}\!\left(\Sigma_{y}+\Sigma_{y\mid x_{j}}-2\big(\Sigma_{y}^{1/2}\Sigma_{y\mid x_{j}}\Sigma_{y}^{1/2}\big)^{1/2}\right)}_{\text{covariance shift}}.$ (18)
@@
Our proposed ICE-based metrics are conceptually related to the Wasserstein–Bures framework in that they distinguish between average effects and variability effects; however, the correspondence is not algebraic in general. In the Wasserstein–Bures decomposition, the covariance shift term is defined through the conditional covariance $\Sigma_{y\mid x_{j}}$, that is, the variability of $y$ across the complementary inputs $\boldsymbol{x}_{C}$ for a fixed $x_{j}$. In contrast, the ICE-based dispersion quantifies the variability of $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$ across different values of $x_{j}$ for a fixed realization $\boldsymbol{x}_{C}^{(i)}$. Thus, the two metrics probe variability along different conditional directions: $\Sigma_{y\mid x_{j}}$ looks across $\boldsymbol{x}_{C}$ at fixed $x_{j}$, whereas ICE-based dispersion looks across $x_{j}$ at fixed $\boldsymbol{x}_{C}^{(i)}$. The ICE-based dispersion should therefore be interpreted as a complementary measure of response heterogeneity rather than a direct estimator of the Wasserstein covariance shift. Under additional structural assumptions (such as additive or separable models, or finite orthonormal expansions like PCE), these perspectives can be related. A direct Wasserstein-based dispersion measure would involve the conditional standard deviation $\sigma_{y\mid x_{j}}=\sqrt{\mathrm{Var}(y\mid x_{j})}$, which quantifies variability across $\boldsymbol{x}_{C}$ for a fixed $x_{j}$, whereas ICE-based dispersion captures the variability of $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$ across different values of $x_{j}$ for a fixed realization $\boldsymbol{x}_{C}^{(i)}$.
Note that the variability captured by $I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)$ will always account for more localized variations specific to $\boldsymbol{x}_{C}^{(i)}$, which are not captured by the globally averaged $\hat{f}_{pdp}(x_{j})$. As a result, we demonstrate, in a certain setting, that:
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j}).$ (19)
More precisely, we demonstrate that the inequality holds for several classes of functions, most notably for those expressible as a finite expansion of a linear combination of bases and functions with some combination of additive and multiplicative separability, as detailed in Appendix A. Since the PCE surrogate model itself constitutes a finite expansion of such bases, the inequality is satisfied when PCE is employed.
The inequality in Eq. (19) may appear analogous to the well-known inequality in the Sobol’ index, namely that the total Sobol’ index is always greater than or equal to the first-order Sobol’ index. However, the two results arise from fundamentally different mechanisms. The Sobol’ inequality follows directly from the variance-based decomposition, whereby the total index aggregates the main effect and all interaction effects in an additive manner. In contrast, the ICE-based inequality reflects the fact that averaging conditional response variations preserves heterogeneous effects that may cancel out when the model response is averaged first, as in the PDP.
@@ 4.2 ICE-based correlation values
Beyond quantifying GSA, it is also essential to assess how interactions modify the relationship between inputs and predictions. This enables the identification of variables that change from positive to negative associations when interacting with others. To complement the ICE-based feature importance, we introduce the ICE-based correlation values which quantify the correlation between ICE curves and the partial dependence function:
@@
where Corr denotes the correlation. That is, we compute the correlation between the ICE curve $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$ for a specific instance $\boldsymbol{x}_{C}^{(i)}$ and the corresponding partial dependence function $\hat{f}_{pdp}(x_{j})$ for the variable $x_{j}$. In this paper, we measure the linear relationship between ICE curves and PDP using the Pearson correlation coefficient to ease interpretation. Alternative methods, including Spearman correlation, could be used to detect monotonic dependencies.
@@
which is essentially the standard deviation of the ICE correlation values, and can be computed from a Monte Carlo simulation. If $\sigma_{\rho}=0$, this indicates that the partial dependence function is similar to ICE curves (in other words, there is no interaction). Conversely, a high value of $\sigma_{\rho}$ suggests that most of the ICE curves differ significantly from the partial dependence line, indicating a strong interaction effect on the trend. Additionally, showing the boxplot or violin plot of ICE-based correlation values also helps in the assessment. It is worth noting that, although Pearson correlation is used, the metric $\sigma_{\rho_{x_{j}}}$ is capable of detecting nonlinear interactions in terms of trend shifts, as it quantifies the variability in correlation between the PDP and multiple ICE curves.
The numerical procedure is shown in Algorithm 2. The computation of the ICE-based correlation metric in Eq. (20) requires evaluating the Pearson correlation between an ICE curve and the corresponding PDP, both discretised on a grid of $K$ points. Computing a single correlation therefore scales as $\mathcal{O}(K)$. Repeating this procedure across $N$ realizations of the complementary inputs $\boldsymbol{x}_{C}$ results in an overall complexity of $\mathcal{O}(NK)$ per input variable. The subsequent computation of the mean and standard deviation in Eq. (21) incurs only $\mathcal{O}(N)$ additional cost and is negligible in comparison. When applied to all $m$ input input variables, the total computational complexity scales as $\mathcal{O}(mNK)$.
Algorithm 2 ICE-based correlation values for a variable $x_{j}$
Trained model $\hat{f}(\boldsymbol{x})$, input variable of interest $x_{j}$, $K$ grid points $\{x_{j}^{(k)}\}_{k=1}^{K}$, partial dependence function $\hat{f}_{pdp}(x_{j})$
$\sigma_{\rho_{x_{j}}}$
Sample $\{\boldsymbol{x}_{C}^{(i)}\}_{i=1}^{N}$ from $\boldsymbol{\pi}_{\boldsymbol{x}_{C}}$
for $i=1,\dots,N$ do
  Evaluate the $i$-th ICE curve $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$
  Compute the correlation between $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$ and $\hat{f}_{pdp}(x_{j})$, i.e., $\rho(x_{j},{\boldsymbol{x}_{C}^{(i)}})$ from Eq. (20)
end for
Compute the standard deviation of ICE correlation values $\sigma_{\rho_{x_{j}}}$ from Eq. (21)
It is important to note that the assessment of the ICE-based correlation value is meaningful only when the partial dependence function exhibits non-zero variance. In cases where the PDP is flat, indicating the absence of a main effect of the corresponding input variable, the Pearson correlation is mathematically undefined due to zero variance. However, in practical applications, a PDP that is theoretically flat may exhibit small but non-zero variance due to the numerical averaging of heterogeneous ICE curves or surrogate approximation errors. To avoid spurious interpretations, the proposed ICE-based correlation metrics are interpreted only when the following criterion is satisfied:
$\frac{I_{pdp}}{\mu_{I_{ice}}}<\varepsilon,$ (22)
where $\varepsilon$ is the threshold criterion. Since $I_{pdp}$ quantifies the importance according to PDP, while $\mu_{I_{ice}}$ captures the average conditional variability across ICE curves, this criterion identifies interaction-dominated variables even when small numerical PDP variations are present. In such cases, ICE-based correlation metrics are not interpreted. Regardless, when this condition is satisfied, it indicates a substantial shift in the conditional trends by which the input variable influences the output, warranting a closer examination of the corresponding ICE curves. In this paper, we use $\varepsilon=5\times 10^{-3}$, such that the PDP-based importance must be less than 0.5% of the average ICE-based importance.
4.3 Evaluation workflow
In practice, the proposed sensitivity analysis framework is applied through a combination of visual diagnostics and quantitative ICE-based metrics. As a prerequisite, the model response function must first be available, either in the form of the original deterministic model or through a sufficiently accurate surrogate model. This requirement is essential, since sensitivity measures derived from a surrogate that does not adequately reproduce the underlying system behavior may be misleading and unreliable [40].
In this paper, we use the coefficient of determination, $R^{2}$, to assess the accuracy of the surrogate model:
$R^{2}=1-\frac{\sum_{i=1}^{n_{v}}\left(f(\boldsymbol{x}^{(i)})-\hat{f}(\boldsymbol{x}^{(i)})\right)^{2}}{\sum_{i=1}^{n_{v}}\left(f(\boldsymbol{x}^{(i)})-\bar{f}\right)^{2}},$ (23)
where $\bar{f}=\frac{1}{n_{v}}\sum_{i=1}^{n_{v}}f(\boldsymbol{x}^{(i)})$ and $n_{v}$ is the size of validation samples. The use of a surrogate model introduces additional uncertainty into the computed GSA metrics due to surrogate approximation error. Quantifying and propagating this uncertainty is beyond the scope of the present work and will be addressed in future studies. In the present analysis, we therefore require the surrogate model to be sufficiently accurate so that the extracted sensitivity information remains faithful to the underlying response behavior. To this end, we require the $R^{2}$ to exceed 0.95 to ensure a faithful representation of the extracted knowledge [36].
The analysis begins with visualization by inspecting PDPs and the corresponding ICE curves to obtain an initial qualitative understanding of global trends and conditional behavior. Subsequently, the ICE-based importance metric $\mu_{I_{\mathrm{ice}},x_{j}}$ is computed to assess the overall importance of each input variable in terms of the average magnitude of its conditional effect. The variability metric $\sigma^{2}_{I_{\mathrm{ice}},x_{j}}$ is then analyzed to quantify heterogeneity across conditional settings; a large value of $\sigma^{2}_{I_{\mathrm{ice}},x_{j}}$ indicates that the influence of $x_{j}$ strongly depends on the remaining inputs and suggests the presence of interaction effects. In such cases, the ICE curves are revisited to examine how the interaction manifests across different realizations of the complementary variables. Finally, the correlation between individual ICE curves and the corresponding PDP is evaluated to assess whether the average PDP is representative of the conditional response behavior.
It should be noted that $\sigma^{2}_{I_{ice}}$ does not directly identify which input variables an input $x_{j}$ strongly interacts with. Such an interaction structure can be explored qualitatively through visualisation, for example, by plotting ICE curves coloured by another input variable, where clear trends indicate potential interactions. However, these visual diagnostics do not provide a quantitative measure of interaction magnitude. In such cases, variance-based interaction metrics, such as second-order Sobol’ indices, can be employed to quantify pairwise interaction effects explicitly. Subsequently, the ICE curves are colour-coded according to the variable with which strong interactions are observed. Therefore, the proposed ICE-based metrics are complementary to other metrics, such as Sobol’ indices, in which the former focuses on behavioral insight from conditional response analysis.
We emphasize that the primary focus of this work is the development of sensitivity metrics themselves, rather than uncertainty quantification of surrogate models. In this context, surrogate models are employed as computational enablers to emulate expensive simulations and to facilitate the application of the proposed sensitivity analysis framework. As such, the proposed metrics are applicable to any function or a surrogate model. When surrogate models are used, additional sources of uncertainty associated with model approximation and training data inevitably arise. While quantifying and propagating these surrogate-induced uncertainties is an important topic, it is beyond the scope of the present study. In this work, we ensure that sufficiently accurate surrogate models are available, allowing us to focus on extracting reliable sensitivity insights from the learned response. In practical settings, surrogate uncertainty can be assessed using established techniques such as bootstrapping, which can be naturally integrated with the proposed framework in future extensions.